# Optimal multipole center for subwavelength acoustic scatterers

### v2

@@ Abstract
Figure 1: Sketch of the considered problem. (a) A cone of height $h$ under insonification of an acoustic plane wave propagating along the $z$-axis. The center of the multipole expansion (green point) is offset from the center of mass (pink point $O$, $z=0$) by a distance $d$. (b) First three multipole moments of the cone vs. $d$ at a fixed frequency. The quadrupole moment exhibits a minimum, whereas the sum of the monopole and dipole moments has a maximum, manifesting the optimal multipole center’s position $z=d_{\rm opt}$.
@@
We begin with a mathematical formulation of our problem. The scattered pressure $p_{\rm sca}(k;\mathbf{r})$ of an acoustical scatterer can be expanded into a truncated expansion of scalar spherical waves (multipoles) weighted with multipole moments Blackstock (2000)
@@
$\ell$ and $m$ is the order and degree of the singular multipole $\psi^{(3)}_{\ell,m}(k;\mathbf{r}-\mathbf{d})$ [see Sec. S1A in Supplementary material (SM) for its definition]. $\mathbf{r}\neq\mathbf{d}$, and $k=\omega/c_{0}$ is the wavenumber in the surrounding medium. $\mathbf{d}$ is an offset of the multipole center relative to the scatterer’s center of mass defined as $\mathbf{r}_{\rm CM}=\bm{0}$ [point $O$ in Fig. 1(a)]. For Eq. (1) to be valid, the coordinate $\mathbf{r}$ must belong to the space region outside a sphere circumscribing the scatterer. Auguié et al. (2016) We always want to choose the optimal $\mathbf{d}_{\rm opt}$ that simultaneously provides the best precision for calculating the physical quantities and the lowest maximal multipole order $\ell_{\rm max}$ in the expansion. In this work, we assume that the offset $\mathbf{d}$ is subwavelength, that is, $k|\mathbf{d}|\ll 1$.
@@
Although standard numerical methods Chong and Żak (2013) can be used to find $\mathbf{d}_{\rm opt}$, we solve this problem analytically. We start by adapting (2) to the considered case and write the off-origin quadrupole moment as a function of a normalized offset $(kd)$. For all zonal multipole coefficients, we use substitutions $p_{2}(kd)\equiv p_{2,0}(k;d\hat{\mathbf{z}})$, and $p_{\ell}\equiv p_{\ell,0}(k;\bm{0})$ for brevity, and then arrive at the following compact equation
@@
where $f_{0,k}=|p_{2}|^{2}$, $f_{1,k}=-\frac{4}{\sqrt{15}}\Re\left[p_{1}^{*}p_{2}\right]$, and $f_{2,k}=\frac{4}{15}|p_{1}|^{2}-\frac{11}{21}|p_{2}|^{2}+\frac{2}{3\sqrt{5}}\Re\left[p_{0}^{*}p_{2}\right]$ with a real $k$ [for a complex $k$, we refer to Sec. S2C in SM] and $p_{\ell}=p_{\ell}(k)$, calculated for $d=0$. The accuracy of Eqs. (3) and (5) in the considered spectral range is confirmed in Sec. S2A in SM. Similar interpolations can be derived for the coefficients $p_{0}(kd)$ and $p_{1}(kd)$ [see Sec. S2B in SM] as well as for high-order terms. The main outcome of (5) is an analytical formula for the optimal offset being a solution to the equation $\partial_{(kd)}|\tilde{p}_{2}(kd)|^{2}=0$ ($k$ is fixed) that provides $d_{\rm opt}=-f_{1,k}/\left(2kf_{2,k}\right)$, or
@@
where the sign of the second derivative $f_{2,k}>0$ and $k\neq 0$ have been implied. Thus, once the multipole moments at $d=0$ are known, the OMC’s position is directly given by (6). Moreover,  (6) can be used even if the initial multipole center is at an arbitrary point $d=d_{0}$. Then, a translation $\left(d_{\rm opt}+d_{0}\right)$ gives the OMC.
@@
Figure 2: (a) Scattering efficiency $\sigma_{\rm tot}(k)/(0.25\pi h^{2})$ of the cone, shown in the inset, with its multipole decomposition for the multipole center at $d=0$. (b) Normalized zonal quadrupole moment of the cone $|p_{2}(k;d)|^{2}/\sum_{\ell\neq 2}|p_{\ell}(k;d)|^{2}$ for $\ell_{\rm max}=5$. The green dots show the values of $d_{\rm opt}=\min_{d}|p_{2}(k;d)|^{2}$, obtained directly from the full-wave numerical simulations. The dashed cyan and solid red lines show $d_{\rm opt}$ calculated with the quadratic [Eq. (6)] and quartic [Eq.  (8)] approximations of $|p_{2}(k;d)|^{2}$, respectively. (c) Error of the scattering cross-section for $\ell_{\rm max}=1$ relative to that for $\ell_{\rm max}=5$ at $d=0$ (pink dashed) and $d=d_{\rm opt}$, calculated in the full-wave model (green solid). Inset: the same but on a logarithmic scale for $\omega h/c_{0}\leq 2.5$.
To validate the analytical results, we perform a numerical simulation of the scattering of a plane wave by a cone of height $h=1$ cm and a radius of $h/2$. The cone is made of a material with the speed of sound $c_{1}$ and mass density $\rho_{1}$. It is placed in a medium with $c_{0}/c_{1}=\sqrt{21}$ and $\rho_{0}/\rho_{1}=1/7$ [see Fig. 1(a)]. We use our in-house T-matrix-based code acoustotreamsaco which is an adaptation of the electromagnetic scattering code treams for acoustic scattering. Beutel, Fernandez-Corbaton, and Rockstuhl (2024) The acoustic T-matrix of the cone was calculated for $d=0$ in a spectral range between 5-25.5 kHz with a step of 0.1 kHz using the finite-element method (Pressure Acoustics interface in COMSOL Multiphysics, v. 6.2). com The acoustic T-matrix was then exported to acoustotreams to calculate the offset multipole moments (2) with $\ell_{\rm max}=5$ (full-wave model).
@@
First, the OMC converges to $d_{\rm opt}=-0.075$ cm in a subwavelength regime $\omega h/c_{0}\lesssim 1$ [see Fig. 2(b)]. Thus, the OMC is located below the center of mass, similar to the magnetic multipole center of electromagnetic scatterers. Kildishev, Achouri, and Smirnova (2024) The OMCs predicted by all three models coincide for $\omega h/c_{0}<2.75$ except at the monopole resonance $\omega_{\rm I}h/c_{0}=1.68$. At this frequency, the quartic approximation $|\breve{p}_{2}|^{2}$ fails near the actual minimum of $|p_{2}|^{2}$ taking nonphysical negative values. Therefore, the sextic approximation should be used. We do not show it here for brevity, but it can be found in Sec. (S2.C.4) in SM. For optical scatterers, we have already demonstrated that the optimal multipole center differs from the scatterer’s center of mass and is, in fact, dispersive Kildishev, Achouri, and Smirnova (2024). In contrast, at dipole resonance $\omega_{\rm II}h/c_{0}=2.4$, all models predict the same OMC with a significant suppression of the error [see the inset in Fig. 2(c)]. In Fig. 2(b), we also see that the negative OMC turns into a large positive OMC around $\omega_{\rm III}h/c_{0}\simeq 2.75$ (the gray region in Fig. 2), and the second derivatives of (5) and (7) become negative [see Sec. S3 in SM]. In this case, the monopole contribution to the scattering is suppressed, while the quadrupole becomes substantial. Figure 2(c) also confirms that neglecting the quadrupole around $\omega\simeq\omega_{\rm III}$ leads to an error larger than $40$%. Hence, the OMC should be redefined as a minimum of a higher-order multipole moment. Surprisingly, at the quadrupole resonance $\omega_{\rm IV}h/c_{0}=3.24$, the quadrupole moment can still be neglected with an error of $10$% since the monopole also exhibits a resonant-like response. Thus, a center for the multipole expansion of the fields placed at the optimal point provides the best conditions for MDA. The error in the scattering cross-section is consistently smaller when the OMC is chosen for the expansion compared to when the center of mass is chosen [see Fig. 2(c)]. At the same time, one can use Eq. (6) to predict the positions of the OMC for $\omega h/c_{0}\leq 2.75$ when $f_{2,k}>0$. For higher frequencies, quartic (8) or the higher approximations should be employed [see Fig. 2(b)].
@@
In a further analysis, we demonstrate that the OMC can be found not only for individual scatterers but also for finite arrangements thereof. The framework of the T-matrix method can be efficiently applied to acoustic multi-scattering problems Peterson and Ström (1975); Kafesaki and Economou (1999); Psarobas, Stefanou, and Modinos (2000), where one needs to distinguish a formulation in a local and global description of the scattered fields of the arrangements Suryadharma et al. (2017). In the local description (LD), the fields are expanded at the $N$ local positions of all $N$ scatterers, $\mathbf{r}_{i}$ with $i=1..N$. The multiple scattering equation is solved to obtain the self-consistent multipole moments that account for the interactions of the scatterers with the incident field and with each other (see, e.g., (17) in Ref. [Beutel, Fernandez-Corbaton, and Rockstuhl, 2024]). In the global description (GD), the fields are expanded from all particles using a global expansion center at a single point $\mathbf{d}$
@@
Equation (9) is similar to (2) for a single-particle case but with the sum over different particles. The other feature is more inconspicuous. The maximal multipole order in the GD $L_{\rm max}$ can be larger than in the LD $\ell_{\rm max}$. Thus, the transition from the LD to the GD is reasonable when the size of an arrangement T-matrix in the GD, $(L_{\rm max}+1)^{2}$, is smaller than the size of an arrangement T-matrix in the LD, $N(\ell_{\rm max}+1)^{2}$.
As an example, we consider a trimer of spherical particles shown in Fig. 3(a). The spheres of radius $R=0.5$ cm are placed at the nodes of an equilateral triangle of side $a=3R$. The material parameters (mass density and speed of sound) of the scatterer material and background are the same as before. Hereafter, we can use the same formalism as before but in a multi-scattering scenario to find the minima of the quadrupole moment in the GD $\sum_{m=-2}^{2}|p^{\rm(global)}_{2,m}(k;\mathbf{d})|^{2}$. An interaction between the spheres can induce the multipoles with $m\neq 0$. However, their contribution is smaller than that of zonal multipoles. Therefore, we can apply Eqs. (6) and (8) to predict $d_{\rm opt}$. We assume the range of frequencies 1–11 kHz where $\ell_{\rm max}=1$ and $L_{\rm max}=2$ are sufficient. If we manage to find the OMC where even $L_{\rm max}=1$ provides good accuracy, then we will decrease the size of the trimer T-matrix from 12 to 4, although one has to track $d_{\rm opt}(\omega)$.
@@
Figure 3 presents the same content as Fig. 2 but for the trimer. The main difference lies in the behavior of $d_{\rm opt}$. In the static case, the system is symmetric concerning the rotations by $\pi/3$ and $32\pi/$ 3 in the $xz$ plane; hence $d_{\rm opt}\to 0$ when $\omega a^{\prime}/c_{0}\to 0$, i.e., the OMC coincides with the center of mass [see Fig. 3(b)]. Here, $\omega a^{\prime}/c_{0}$ is an effective trimer size parameter with $a^{\prime}=a+2R$. As $\omega a^{\prime}/c_{0}$ increases, $d_{\rm opt}$ also increases. From $\omega a^{\prime}/c_{0}=2.6$, the OMC remains at the same point $d_{\rm opt}\approx 0.35$ cm close to the center of the upper sphere and provides a better accuracy of the MDA [see Fig. 3(c)]. Both quadratic and quartic models predict $d_{\rm opt}$ for $\omega a^{\prime}/c_{0}\leq 2.6$ very well, although the quadratic model fails for $\omega a^{\prime}/c_{0}>2.6$ due to the resonant response of the trimer [see Fig. 3(a)]. In Sec. S4 in SM, we also consider this trimer rotated by $\pi/2$ in the $xz$ plane, for which $\mathbf{d}_{\rm opt}$ is traced as a trajectory, with the initial point at $\mathbf{d}_{\rm opt}=\bm{0}$, as the frequency increases.
@@
In the end, we demonstrate that the OMC can improve the computational efficiency of the acoustic T-matrix method in multiple scattering problems. For that purpose, we consider a ring of the $N$ identical cones (the same as those previously considered) placed at a distance $D=2.6$ cm between the nearest neighbors. The acoustic T-matrix of the individual cone is calculated at frequency $\omega h/c_{0}=2.5$ for the multipole expansion center placed at $d=0$ with $\ell_{\rm max}=2$, and at the OMC with $\ell_{\rm max}=1$ [see Fig. 2(b)]. An additional modification of the OMCs due to the interaction between scatterers is neglected. Then, we vary $N$ from 80 to 380, estimate the time we need to solve the multiple scattering problem, compute the ring scattering cross section for each $N$, average this time after 10 iterations, and normalize it by $\tau$. Here, $\tau$ is the average time to solve a linear system with a $1000\times 1000$ matrix and $1000\times 1$ right-hand side using the numpy.linalg.solve function in Python 3.11.
Figure 4 illustrates the results, demonstrating that the OMC reduces computational time by six since we manage to reduce the size of the isolated-cone T matrix from 9 for $\ell_{\rm max}=2$ to 4 for $\ell_{\rm max}=1$, while the error in the calculation of the ring scattering cross-section remains comparable. Moreover, using the MDA but placing the multipole center at the center of mass increases the average error from 1.42% to 4%. Thus, the OMC remarkably reduces the computational time within the T-matrix and multipole methods for simulating the acoustic response of discrete scatterers while maintaining a reasonable error.
@@
In conclusion, we address the problem of determining the optimal point to place the center of the multipole expansion in acoustic scattering. We consider individual scatterers and multiple-particle structures without reflection symmetry along the $z$-axis and trace the OMC position along the axis upon insonification of the scatterers with a plane wave. Our comprehensive acoustic scattering model reveals that similar to electromagnetic scattering, Kildishev, Achouri, and Smirnova (2024) an optimal solution that minimizes the multipole content to produce the most efficient descriptor while retaining only a few terms of the lowest degree. We showcase the advantages of using low-degree multipoles (monopole and dipole) at this OMC, demonstrating substantial computational efficiency gains achieved in T-matrix calculations. Our approach can be extended beyond the systems described with the monopole-dipole-quadrupole approximation to systems of arbitrary geometry that require optimizing their multipole descriptions up to arbitrary degrees and orders.
See the Supplementary material for the definition of scalar spherical waves (multipoles), the addition theorem and translation coefficients, the derivation of Eqs. (3)-(8), the analysis of the error of the approximations (3),  (4), and (7), calculation of an optimal multipole center trajectory for the rotated trimer, the calculation of an optimal multipole center for the isolated cone as a function of the angle of incidence, and the details on the computation of the acoustic T-matrices of axisymmetric objects in COMSOL Multiphysics.

### v3

@@ Abstract
Figure 1: Sketch of the considered problem. (a) A cone of height $h$ under insonification of an acoustic plane wave propagating along the $z$-axis. The center of the multipole expansion (green point) is offset from the center of mass (pink point $O$, $z=0$) by a distance $d$. (b) Squared amplitudes of first three multipole moments of the cone vs. $d$ at a fixed frequency. The quadrupole moment exhibits a minimum, whereas the sum of the monopole and dipole moments has a maximum, manifesting the optimal multipole center’s position $z=d_{\rm opt}$.
@@
We begin with a mathematical formulation of our problem. The scattered pressure $p_{\rm sca}(k;\mathbf{r})$ of an acoustic scatterer can be expressed as a truncated expansion in scalar spherical waves (multipoles) weighted by multipole moments Blackstock (2000)
@@
$\ell$ and $m$ is the degree and order of the singular multipole $\psi^{(3)}_{\ell,m}(k;\mathbf{r}-\mathbf{d})$ [see Sec. S1A in Supplementary material (SM) for its definition]. $\mathbf{r}\neq\mathbf{d}$, and $k=\omega/c_{0}$ is the wavenumber in the surrounding medium. $\mathbf{d}$ is an offset of the multipole center relative to the scatterer’s center of mass defined as $\mathbf{r}_{\rm CM}=\bm{0}$ [point $O$ in Fig. 1(a)]. For Eq. (1) to be valid, the coordinate $\mathbf{r}$ must belong to the space region outside a sphere circumscribing the scatterer. Auguié et al. (2016) We aim to choose the optimal $\mathbf{d}_{\rm opt}$ that simultaneously provides the best precision in calculating physical quantities and the lowest maximal multipole degree $\ell_{\rm max}$ in the expansion. In this work, we assume that the offset $\mathbf{d}$ is subwavelength, that is, $k|\mathbf{d}|\ll 1$.
@@
Although standard numerical methods Chong and Żak (2013) can be used to find $\mathbf{d}_{\rm opt}$, we solve this problem analytically. We start by adapting (2) to the considered case and write the off-origin quadrupole moment as a function of a normalized offset $(kd)$. For all zonal multipole coefficients, we make substitutions $p_{2}(kd)\equiv p_{2,0}(k;d\hat{\mathbf{z}})$, and $p_{\ell}\equiv p_{\ell,0}(k;\bm{0})$ for brevity, and then arrive at the following compact equation
@@
where $f_{0,k}=|p_{2}|^{2}$, $f_{1,k}=-\frac{4}{\sqrt{15}}\Re\left[p_{1}^{*}p_{2}\right]$, and $f_{2,k}=\frac{4}{15}|p_{1}|^{2}-\frac{11}{21}|p_{2}|^{2}+\frac{2}{3\sqrt{5}}\Re\left[p_{0}^{*}p_{2}\right]$ with a real $k$ [for a complex $k$, we refer to Sec. S2C in SM] and $p_{\ell}=p_{\ell}(k)$, calculated for $d=0$. The accuracy of Eqs. (3) and (5) in the considered spectral range is confirmed in Sec. S2A in SM. Similar interpolations can be derived for the coefficients $p_{0}(kd)$ and $p_{1}(kd)$ [see Sec. S2B in SM] as well as for high-degree terms. The main outcome of (5) is an analytical formula for the optimal offset being a solution to the equation $\partial_{(kd)}|\tilde{p}_{2}(kd)|^{2}=0$ ($k$ is fixed) that provides $d_{\rm opt}=-f_{1,k}/\left(2kf_{2,k}\right)$, or
@@
where the sign of second derivative $f_{2,k}>0$ and $k\neq 0$ have been implied. Thus, once the multipole moments at $d=0$ are known, the OMC’s position is directly given by (6). Moreover,  (6) can be used even if the initial multipole center is at an arbitrary point $d=d_{0}$. Then, a translation $\left(d_{\rm opt}+d_{0}\right)$ gives the OMC.
@@
Figure 2: (a) Scattering efficiency $\sigma_{\rm tot}(k)/(0.25\pi h^{2})$ of the cone, shown in the inset, with its multipole decomposition for the multipole center at $d=0$. (b) Normalized zonal quadrupole moment of the cone $|p_{2}(k;d)|^{2}/\sum_{\ell\neq 2}|p_{\ell}(k;d)|^{2}$ for $\ell_{\rm max}=5$. The green dots show the values of $d_{\rm opt}=\min_{d}|p_{2}(k;d)|^{2}$, obtained directly from the full-wave numerical simulations. The dashed blue and solid red lines show $d_{\rm opt}$ calculated with the quadratic [Eq. (6)] and quartic [Eq.  (8)] approximations of $|p_{2}(k;d)|^{2}$, respectively. (c) Error of the scattering cross-section for $\ell_{\rm max}=1$ relative to that for $\ell_{\rm max}=5$ at $d=0$ (pink dashed) and $d=d_{\rm opt}$, calculated in the full-wave model (green solid). Inset: the same but on a logarithmic scale for $\omega h/c_{0}\leq 2.5$.
To validate the analytical results, we perform a numerical simulation of the scattering of a plane wave by a cone of height $h=1$ cm and radius of $h/2$. The cone is made of a material with the speed of sound $c_{1}$ and mass density $\rho_{1}$. It is placed in a medium with $c_{0}/c_{1}=\sqrt{21}$ and $\rho_{0}/\rho_{1}=1/7$ [see Fig. 1(a)]. We use our in-house T-matrix-based code acoustotreamsaco which is an adaptation of the electromagnetic scattering code treams for acoustic scattering. Beutel, Fernandez-Corbaton, and Rockstuhl (2024) The acoustic T-matrix of the cone was calculated for $d=0$ in a spectral range between 5-25.5 kHz with a step of 0.1 kHz using the finite-element method (Pressure Acoustics interface in COMSOL Multiphysics, v. 6.2). com The acoustic T-matrix was then exported to acoustotreams to calculate the offset multipole moments (2) with $\ell_{\rm max}=5$ (full-wave model).
@@
First, the OMC converges to $d_{\rm opt}=-0.075$ cm in a subwavelength regime $\omega h/c_{0}\lesssim 1$ [see Fig. 2(b)]. Thus, the OMC is located below the center of mass, similar to the magnetic multipole center of electromagnetic scatterers. Kildishev, Achouri, and Smirnova (2024) The OMCs predicted by all three models coincide for $\omega h/c_{0}<2.75$ except at the monopole resonance $\omega_{\rm I}h/c_{0}=1.68$. At this frequency, the quartic approximation $|\breve{p}_{2}|^{2}$ fails near the actual minimum of $|p_{2}|^{2}$ taking nonphysical negative values. Therefore, the sextic approximation should be used. We do not show it here for brevity, but it can be found in Sec. (S2.C.4) in SM. For optical scatterers, we have already demonstrated that the optimal multipole center differs from the scatterer’s center of mass and is, in fact, dispersive Kildishev, Achouri, and Smirnova (2024). In contrast, at dipole resonance $\omega_{\rm II}h/c_{0}=2.4$, all models predict the same OMC with a significant suppression of the error [see the inset in Fig. 2(c)]. In Fig. 2(b), we also see that the negative OMC turns into a large positive OMC around $\omega_{\rm III}h/c_{0}\simeq 2.75$ (the gray region in Fig. 2), and the second derivatives of (5) and (7) become negative [see Sec. S3 in SM]. In this case, the monopole contribution to the scattering is suppressed, while the quadrupole becomes substantial. Figure 2(c) also confirms that neglecting the quadrupole around $\omega\simeq\omega_{\rm III}$ leads to an error larger than $40$%. Hence, the OMC should be redefined as a minimum of a higher-degree multipole moment. Surprisingly, at the quadrupole resonance $\omega_{\rm IV}h/c_{0}=3.24$, the quadrupole moment can still be neglected with an error of $10$% since the monopole also exhibits a resonant-like response. Thus, a center for the multipole expansion of the fields placed at the optimal point provides the best conditions for the MDA. The error in the scattering cross-section is consistently smaller when the OMC is chosen for the expansion compared to when the center of mass is chosen [see Fig. 2(c)]. At the same time, one can use Eq. (6) to predict the positions of the OMC for $\omega h/c_{0}\leq 2.75$ when $f_{2,k}>0$. For higher frequencies, Eq. (8) or the higher approximations should be employed [see Fig. 2(b)].
@@
In a further analysis, we demonstrate that the OMC can be found not only for individual scatterers but also for finite arrangements thereof. The framework of the T-matrix method can be efficiently applied to acoustic multiple scattering problems Peterson and Ström (1975); Kafesaki and Economou (1999); Psarobas, Stefanou, and Modinos (2000), where one needs to distinguish a formulation in a local and global description of the scattered fields of arrangements Suryadharma et al. (2017). In the local description (LD), the fields are expanded at the $N$ local positions of all $N$ scatterers, $\mathbf{r}_{i}$ with $i=1..N$. The multiple scattering equation is solved to obtain the self-consistent multipole moments that account for the interactions of the scatterers with the incident field and with each other (see, e.g., (17) in Ref. [Beutel, Fernandez-Corbaton, and Rockstuhl, 2024]). In the LD, the multipole moments are calculated when the multipole centers coincide with the centers of mass of scatterers located at $\mathbf{r}_{i}$. In the global description (GD), the fields are expanded from all particles using a global expansion center at a single point $\mathbf{d}$. In this regard, the global multipole moments depend on $\mathbf{d}$ and the local multipole moments as
@@
Equation (9) is similar to (2) for the single-particle case, but it includes a sum over the scatterers. The other feature is more inconspicuous. The maximal multipole degree in the GD $L_{\rm max}$ can be larger than in the LD $\ell_{\rm max}$. Thus, the transition from the LD to the GD is reasonable when the size of an arrangement T-matrix in the GD, $(L_{\rm max}+1)^{2}$, is smaller than the size of an arrangement T-matrix in the LD, $N(\ell_{\rm max}+1)^{2}$.
As an example, we consider a trimer of spherical particles shown in Fig. 3(a). The spheres of radius $R=0.5$ cm are placed at the nodes of an equilateral triangle of side $a=3R$. The material parameters (mass density and speed of sound) of the scatterer material and background are the same as before. Hereafter, we can use the same formalism as before but in a multiple scattering scenario to find the minima of the quadrupole moment in the GD $\sum_{m=-2}^{2}|p^{\rm(global)}_{2,m}(k;\mathbf{d})|^{2}$. An interaction between the spheres can induce the multipoles with $m\neq 0$. However, their contribution is smaller than that of zonal multipoles. Therefore, we can apply Eqs. (6) and (8) to predict $d_{\rm opt}$. We assume the range of frequencies 1–11 kHz where $\ell_{\rm max}=1$ and $L_{\rm max}=2$ are sufficient. If we manage to find the OMC where even $L_{\rm max}=1$ provides good accuracy, then we will decrease the size of the trimer T-matrix from 12 to 4, although one has to track $d_{\rm opt}(\omega)$.
@@
Figure 3 presents the same content as Fig. 2 but for the trimer. The main difference lies in the behavior of $d_{\rm opt}$. In the static case, the system is symmetric concerning the rotations by $\pi/3$ and $32\pi/$ 3 in the $xz$ plane; hence $d_{\rm opt}\to 0$ when $\omega a^{\prime}/c_{0}\to 0$, i.e., the OMC coincides with the center of mass of the trimer [see Fig. 3(b)]. Here, $\omega a^{\prime}/c_{0}$ is an effective trimer size parameter with $a^{\prime}=a+2R$. As $\omega a^{\prime}/c_{0}$ increases, $d_{\rm opt}$ also increases. From $\omega a^{\prime}/c_{0}=2.6$, the OMC remains at the same point $d_{\rm opt}\approx 0.35$ cm close to the center of the upper sphere and provides a better accuracy of the MDA [see Fig. 3(c)]. Both quadratic and quartic models predict $d_{\rm opt}$ for $\omega a^{\prime}/c_{0}\leq 2.6$ very well, although the quadratic model fails for $\omega a^{\prime}/c_{0}>2.6$ due to the resonant response of the trimer [see Fig. 3(a)]. In Sec. S4 in SM, we also consider this trimer rotated by $\pi/2$ in the $xz$ plane, for which $\mathbf{d}_{\rm opt}$ is traced as a trajectory, with the initial point at $\mathbf{d}_{\rm opt}=\bm{0}$, as the frequency increases.
@@
In the end, we demonstrate that the OMC can improve the computational efficiency of the acoustic T-matrix method in multiple scattering problems. For that purpose, we consider a ring of $N$ identical cones (the same as those previously considered) placed at a distance $D=2.6$ cm between the nearest neighbors. The acoustic T-matrix of the individual cone is calculated at frequency $\omega h/c_{0}=2.5$ for the multipole expansion center placed at $d=0$ with $\ell_{\rm max}=2$, and at the OMC with $\ell_{\rm max}=1$ [see Fig. 2(b)]. An additional modification of the OMCs due to the interaction between scatterers is neglected. Then, we vary $N$ from 80 to 380, estimate the time we need to solve the multiple scattering problem, compute the ring scattering cross section for each $N$, average this time after 10 iterations, and normalize it by $\tau$. Here, $\tau$ is the average time to solve a linear system with a $1000\times 1000$ matrix and $1000\times 1$ right-hand side using the numpy.linalg.solve function in Python 3.11.
Figure 4 illustrates the results, demonstrating that the OMC reduces computational time by a factor of six, as we manage to reduce the size of the isolated-cone T-matrix from 9 for $\ell_{\rm max}=2$ to 4 for $\ell_{\rm max}=1$, while the error in the calculation of the ring scattering cross-section remains comparable. Moreover, using the MDA but placing the multipole center at the center of mass increases the average error from 1.42% to 4%. Thus, the OMC remarkably reduces the computational time within the T-matrix and multipole methods for simulating the acoustic response of discrete scatterers while maintaining a reasonable error.
@@
In conclusion, we address the problem of determining the optimal point to place the multipole expansion center in acoustic scattering. We consider individual scatterers and multi-particle structures without reflection symmetry along the $z$-axis and trace the OMC position along the axis when the scatterers are insonified by a plane wave. Our comprehensive acoustic scattering model reveals that similar to electromagnetic scattering, Kildishev, Achouri, and Smirnova (2024) an optimal solution that minimizes the multipole content to produce the most efficient descriptor while retaining only a few terms of the lowest degree. We showcase the advantages of using low-degree multipoles (monopole and dipole) at this OMC, demonstrating substantial computational efficiency gains achieved in T-matrix calculations. Our approach can be extended beyond the systems described within the monopole-dipole-quadrupole approximation to systems of arbitrary geometry that require optimizing their multipole descriptions up to arbitrary multipole degrees and orders.
See the Supplementary material for the definition of scalar spherical waves (multipoles), the addition theorem and translation coefficients, the derivation of Eqs. (3)-(8), the analysis of the error of approximations (3),  (4), and (7), the calculation of an optimal multipole center trajectory for the rotated trimer, the calculation of an optimal multipole center for the isolated cone as a function of the angle of incidence, and the details on the computation of the acoustic T-matrices of axisymmetric objects in COMSOL Multiphysics.